OpenStudy (anonymous):

an object will take t seconds to drop d ft, where t= sqrt d/16. what distance will the object drop in sqrt 13/2 sec?

OpenStudy (anonymous):

okay so \[Time=\sqrt{Distance/16}\] so we have the time now \[\sqrt{13/2}=\sqrt{d/16}\] do you know where to go now?